OpenStudy (anonymous):

Find the equation of the line that is perpendicular to the line 2x+3y-8=0 at the point (1,2)

OpenStudy (anonymous):

first you need to put the equation in y = mx + b form

OpenStudy (anonymous):

y=-2x+8 ----- 3

OpenStudy (anonymous):

2x + 3y - 8 = 0 3y = -2x + 8 y = (-2x + 8) / 3 go a little further... y = -2/3x + 8/3

OpenStudy (anonymous):

understand so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the slope for this equation is - 2/3 but we are looking for the perpendicular slope. That means we are looking for the negative reciprocal of -2/3. All that means is you flip the slope over and change the sign. So the slope we need is 3/2. Do you understand, or have I lost you ?

OpenStudy (anonymous):

no I understand that part but how do I find the rest of the line again by plugging the point I have into the equation?

OpenStudy (anonymous):

use the point slope formula because we have the slope and the points... y - y1 = m(x - x1) using slope(m) = 3/2 and points (1,2) -- x1 = 1 and y1 = 2

OpenStudy (anonymous):

now lets sub them in.... y - 2 = 3/2(x - 1) can you solve it from here ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what did you get ?

OpenStudy (anonymous):

would you like me to finish it and we can compare answers ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

y - 2 = 3/2(x - 1) y - 2 = 3/2x - 3/2 y = 3/2x - 3/2 + 2 y = 3/2x - 3/2 + 4/2 y = 3/2x + 1/2 is this what you got ?

OpenStudy (anonymous):

yes that is what I got thanks

OpenStudy (anonymous):

anytime :)